(h) Sport Alive sold 72 pairs of ski poles. Superlight poles sell at $30 per pair, while ordinary poles sell at $16 per pair. If the total sales value was $1530, how many pairs of each type were sold?

Answer by Boreal(15235)

SL=x
O=y
x+y=72; number equation
30x+16y=1530; cost equation
-30x-30y=-2160
-14y=-630
y=45 O poles ($720)
y=27 SL poles ($810)
Total=$1530